# arXiv:2601.06986v3[physics.ao-ph] 28 Feb 2026

### v3

@@ Abstract
Using a gap-free, millennial-resolution record of atmospheric CO2 and Antarctic temperature since the Last Interglacial (130 ka ~), we structurally identify the feedback architecture of the climate–carbon system. The series are cointegrated, supporting estimation with a Vector Error Correction Model (VECM). The cointegrating relationship implies a temperature response of 13.0 K per CO2 doubling. Structural identification combining a Causality Allocation Curve (CAC) and Milankovitch-cycle instrumental variables yields a contemporaneous carbon response (CCR) of 5.77 ppm/K, while the contemporaneous temperature response (CTR) is statistically indistinguishable from zero. Accounting for lagged feedbacks, the cumulative temperature response within one millennium of CO2 doubling reaches 11.8 K.
@@ Introduction
The statistical analysis of paleoclimate records faces three fundamental challenges. First, the high persistence of temperature and CO2 series carries the risk of spurious regression (Kaufmann et al., 2006; Cummins et al., 2022). Because these variables exhibit the properties of non-stationary, $I(1)$ processes, standard regression techniques can yield misleadingly significant correlations that lack physical causality. Second, a methodological schism exists regarding the appropriate framework for these series; while some advocate for a cointegration approach (Kaufmann et al., 2006; Miller, 2019), others argue that the system is ultimately mean-reverting and stationary, favoring standard Vector Autoregression (VAR) models (Davidson et al., 2016). Third, the irregular temporal spacing of ice-core proxies often necessitates linear interpolation, which has been shown to introduce substantial size distortions and biases in statistical testing (Miller, 2019).
Beyond these statistical hurdles, a long-standing climate-physical debate persists regarding the “lead-lag” relationship between temperature and CO2. Early studies suggested that CO2 lagged Antarctic temperature by several centuries during deglaciations (Caillon et al., 2003), implying that carbon was a follower rather than a driver of initial warming. More recent re-evaluations, accounting for gas-ice age uncertainties, have proposed a nearly synchronous relationship (Parrenin et al., 2013). However, these findings often rely on reduced-form correlations that cannot isolate structural innovations from reciprocal feedbacks. While Lemoine (2010) utilized orbital parameters as instrumental variables (IV) to identify a significantly positive—yet relatively modest—climate–carbon feedback factor, the short-run structural dynamics within a unified millennial-scale framework remain under-explored.
@@
This study addresses these challenges by applying a Structural Vector Error Correction Model (SVECM) to a high-resolution, gap-free ice-core dataset. Our contribution is twofold. First, we identify a specific 130,000-year interval (the 130 ka sample) that is entirely free of observational gaps at a 1,000-year resolution. By utilizing this interval, we eliminate the need for interpolation, thereby physically neutralizing the statistical artifacts warned of by Miller (2019). Second, we employ a structural identification strategy that utilizes the strict exogeneity of Milankovitch cycles as an instrument to disentangle the contemporaneous temperature response (CTR, $a_{12}$) from the contemporaneous carbon response (CCR, $a_{21}$).
@@
Our results reveal a stark asymmetry in the contemporaneous climate–carbon interaction. We find a robust, statistically significant CCR ($\hat{a}_{21}>0$), consistent with the magnitude suggested by last-millennium reconstructions (Frank et al., 2010). In contrast, the CTR ($a_{12}$) is found to be statistically indistinguishable from zero at the millennial scale. This identification implies that while temperature changes trigger an immediate carbon release, the radiative feedback of CO2 onto temperature requires structural lags to manifest. This framework provides a unified explanation for the perceived “synchronicity” in deglacial records as a manifestation of the CCR, while reconciling the “lag” observed in other contexts as the physical adjustment time of the thermal system. By bridging the gap between time-series econometrics and paleoclimatic physics, we provide a more precise characterization of the feedbacks that govern the Earth’s long-term climate sensitivity.
@@
To circumvent this risk while preserving the critical information contained in the levels of the variables, we employ a Vector Error Correction Model (VECM). This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a stable linear combination exists. In this study, we consider a three-variable system $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$, where $T_{t}$ is Antarctic temperature, $\ln C_{t}$ is the logarithm of CO2 concentration, and $M_{t}$ represents external orbital (Milankovitch) forcing. The VECM is expressed as:
$\Delta y_{t}=\alpha\beta^{\prime}y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t}$ (1)
where $\beta$ contains the cointegrating vector, $\alpha$ represents the adjustment speeds to equilibrium, and $\Gamma_{i}$ captures the short-run lagged dynamics. The error term $\epsilon_{t}$ denotes the reduced-form innovations, which are generally correlated.
@@
To isolate structural shocks ($u_{t}$) from the reduced-form residuals ($\epsilon_{t}$), we define the relationship $(I-A_{0})\epsilon_{t}=u_{t}$, where $u_{t}$ is a vector of orthogonal structural innovations with $u_{t}\sim\text{i.i.d.}\mathcal{N}(0,\Sigma_{u})$. In our three-variable system, the contemporaneous interaction matrix $A_{0}$ is defined as:
@@
Equation (2) demonstrates that the allocation of the total contemporaneous correlation between temperature and CO2 into directional causal links is uniquely determined by the sample covariance structure ($s_{ij}$) of the residuals.
@@
The presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline. The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$ (see Appendix A, for details), provides a formal bridge between millennial-scale innovations and the long-term equilibrium states.
@@
To further rigorously test the statistical significance of the CTR ($a_{12}$) while accounting for the uncertainty in the CCR ($a_{21}$), we perform a joint estimation using the Generalized Method of Moments (GMM). This approach incorporates both the regression model in Equation (3) and the orthogonality condition derived from the structural shock properties, known as the Causality Allocation Curve (CAC) or Equation (2). The moment conditions for this joint estimation are defined by the structural identification framework established in Section 2.2.
@@
Figure 3: Identification of structural parameters on the Causality Allocation Curve (CAC). The blue line represents the CAC derived from the residual covariance. The red point and dashed ellipse indicate the GMM joint estimate and the 95% confidence region. The intersection near the $a_{12}=0$ line confirms the unidirectional contemporaneous feedback from temperature to carbon.
@@
Based on the identification results in the previous section, we now evaluate the structural sensitivity of the climate–carbon system. First, to interpret the estimated reduced-form parameters ($\alpha,\beta,\Gamma$) in a structural context, we must consider the transformation $(I-A_{0})$. While the cointegrating vector $\beta$ and the adjustment speeds $\alpha$ remain invariant to this transformation due to the property $\beta^{\prime}\Xi=0$ (as formally derived in Appendix A), the short-run lag coefficients $\Gamma$ must be structurally re-identified using the contemporaneous parameters $a_{12}$ and $a_{21}$.
By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$. The specific element for the CO22-to-temperature lag can be expanded as follows:
@@
This structural consistency allows us to derive the Earth System Sensitivity (ESS) directly from the cointegrating relationship. The estimated long-term equilibrium for the 130 ka sample is given by $T-18.748\ln C=\text{const}$. This relationship implies that a doubling of atmospheric CO2 concentration is associated with a long-term temperature increase of $18.748\times\ln(2)\approx 13.0$ K in Antarctica.
However, since this estimate is based on Antarctic ice core records, it reflects local temperature variations which are subject to polar amplification. Following the methodology of previous studies, we convert this local signal to a Global Average Temperature (GAT) equivalent by applying a scaling factor of approximately 0.5, as Antarctic temperature anomalies are known to be roughly twice as large as the global mean. This conversion yields a global ESS of approximately 6.5 K per CO2 doubling. This value is remarkably consistent with the reconstructed ESS of $9\pm 4.5$ K (95% CI) reported by Snyder (2016), confirming that our VECM-based structural identification captures the fundamental long-term coupling of the Earth’s climate and carbon cycle since the Last Interglacial.
Beyond the long-term equilibrium, the structural VECM allows for a granular decomposition of the transient temperature response. We examine the impact of a doubling of atmospheric CO22 concentration ($\Delta\ln C=\ln(2)$) within a single millennial time step. Under the structural identification a12=0a{12}=0, the immediate temperature response after 1,000 years is driven by two distinct mechanisms: the structural lag effect and the error correction process.
First, the structural lag effect $(\Gamma^{*}_{1})_{12}=15.201$ accounts for a temperature increase of $15.201\times\ln(2)\approx 10.5$ K. This represents the direct millennial-scale sensitivity of the climate system to a carbon shock. Second, the error correction term $\alpha=-0.103$ dictates the speed at which the system reverts to its long-term equilibrium. A doubling of CO2 creates a disequilibrium in the cointegrating relationship of $18.748\times\ln(2)\approx-13.0$ K. The resulting adjustment in the subsequent millennium contributes an additional warming of $(-0.103)\times(-13.0)\approx 1.3$ K.
Summing these effects, the cumulative temperature response within one millennium of a CO2 doubling reaches $11.8$ K in Antarctica. In global average terms, this corresponds to approximately $5.9$ K of warming. While this magnitude appears large, it is remarkably consistent with the transient warming observed in the recent industrial era. For instance, the observed global temperature increase of approximately $1.2$ K over the last 150 years, associated with a CO2 rise from 280 to 428 ppm (a 53% increase), aligns with the sensitivity scale identified in our millennial-resolution model when accounting for the slower response times of the Earth system.
Summing these effects, the cumulative temperature response within one millennium of a CO2 doubling reaches 11.8 K in Antarctica, which translates to a global average warming of approximately 5.9 K. While this 1,000-year response is substantial, it represents only the transient stage of the total adjustment toward the long-term equilibrium (ESS) of 6.5 K.
The proximity of our VECM-based estimates to the results reported by Snyder (2016) warrants further methodological discussion. Snyder (2016) estimated the ESS using a simple OLS framework and was subsequently criticized for potentially ignoring the endogeneity of CO2 and the influence of orbital forcing. However, our identification of a stable cointegrating relationship ($\beta$) provides a strong theoretical justification for the robustness of such OLS-based reconstructions. In the presence of cointegration, OLS estimators possess the property of “super-consistency,” converging to the true long-run parameters even in the face of contemporaneous endogeneity.
By employing a VECM, we have formally verified this long-term stability and separated the structural feedbacks from stochastic innovations. The fact that our structurally identified ESS (6.5 K) and millennial response (5.9 K) align closely with Snyder’s findings suggests that the climate–carbon coupling is so fundamentally dominated by the long-term cointegrating trend that the essence of the Earth system’s sensitivity is captured even by simpler models, provided they span sufficient glacial cycles. Our structural approach thus reinforces the empirical validity of previous reconstructions while providing a rigorous framework to account for internal dynamics and orbital exogeneity.
This millennial sensitivity also provides a critical perspective on modern climate change. A global warming of 5.9 K per CO2 doubling after 1,000 years implies an average linear adjustment rate of approximately 0.885 K per 150 years. While this is numerically lower than the observed 1.2 K increase since the industrial revolution (associated with a 53% rise in CO2), this discrepancy is entirely consistent with the non-linear “front-loading” characteristic of climate responses. The immediate radiative forcing following a carbon shock triggers rapid warming in the upper ocean and atmosphere (often termed the Transient Climate Response), whereas our 6.5 K ESS estimate captures the slow, “committed” warming of the deep Earth system that unfolds over subsequent centuries. Rather than a contradiction, the gap between the modern 1.2 K response and our 1,000-year 5.9 K projection quantifies the massive, yet-to-be-realized thermal debt inherent in the Earth system’s long-term feedbacks.
Finally, we evaluate our estimate of the Contemporaneous Carbon Response (CCR). Our GMM estimation yields $\hat{a}_{21}=0.0206$, which corresponds to a carbon cycle sensitivity of approximately $5.77$ ppm per 1 K of Antarctic warming. This result is highly consistent with the median estimate of $7.7$ ppm/K (likely range: 1.7–21.4 ppm/K) reported by Frank et al. (2010) for the last millennium. The stability of the CCR across different timescales—from the centennial resolution of the late Holocene to the millennial resolution of our 130 ka sample—underscores the robustness of the climate–carbon feedback architecture identified in this study.
@@
Figure 4 illustrates the response to the CO2 doubling shock. Consistent with our null-CTR identification ($a_{12}=0$), temperature shows no immediate response within the same 1,000-year step. Instead, the warming is triggered by structural lag effects, gradually converging toward the long-term ESS of 13.0 K in Antarctica. This dynamic visualization aligns with Rohling et al. (2012), who emphasize that climate sensitivity is a function of the timescale, with the full magnitude of slow feedbacks becoming apparent only as the system approaches a new steady state.
@@
Appendix: Theoretical Background of VECM and Impulse Response
@@
Assuming a constant shock has been applied from the past to the present ($\epsilon_{t-h}=\epsilon$), Equation (6) yields $y_{t}=(\sum_{h=0}^{\infty}\Psi_{h})\epsilon$. On the other hand, assuming a long-run equilibrium state ($y_{t}=y_{t-1}=\dots=y$) and substituting this into Equation (5), we obtain $\left[I-\sum_{i=1}^{p}\Phi_{i}\right]y=Gy=\epsilon$. By comparing the two, the following relationship is derived:
@@
where $C_{j}=\Psi_{j}-\Psi_{j-1}$ represents the impact of the shock $\epsilon_{t-j}$ on the current "change" $\Delta y_{t}$. Given that the level $y_{t}$ is the cumulative sum of past differences $y_{t}=y_{0}+\sum_{j=1}^{t}\Delta y_{j}$, the impulse response for the level, $\Psi_{h}=C_{h}+\Psi_{h-1}$, can be rewritten as:
@@
$\Psi_{h}=\sum_{j=0}^{h}C_{j}=\left(\sum_{j=0}^{\infty}C_{j}\right)-\left(\sum_{j=h+1}^{\infty}C_{j}\right)=\Xi+\Upsilon_{h}$
@@
where $\Pi=\alpha\beta^{\prime}$ represents the long-run equilibrium (cointegrating) relationships. Note that $\Phi_{1}=I+\Pi+\Gamma_{1}$, $\Phi_{i}=\Gamma_{i}-\Gamma_{i-1}$, and $\Phi_{p}=-\Gamma_{p-1}$. Engle and Granger (1987) demonstrated that for $I(1)$ series, a linear combination $\beta^{\prime}y_{t}$ exists that is stationary ($I(0)$).
@@
This mathematical property carries a profound physical meaning: while an exogenous shock (such as a sudden change in orbital forcing or a carbon cycle innovation) can permanently shift the absolute levels of temperature and CO2, it is fundamentally incapable of breaking the equilibrium "bond" between them. The variables move to a new climatic baseline, but the underlying physical relationship, defined by $\beta$, remains intact.

### v4

# Structural Cointegration of the Climate–Carbon Feedback: Evidence from the Last 130,000 Years

@@ Abstract
Using a gap-free, millennial-resolution ice-core record spanning the last 130,000 years, we identify the feedback architecture between Antarctic temperature and atmospheric CO2. The series are found to be cointegrated, justifying estimation with a Vector Error Correction Model (VECM). The estimated long-run relationship yields a temperature change of 13.0 K per CO2 doubling. Structural identification combining Milankovitch-cycle instrumental variables with the VECM residuals yields a contemporaneous carbon response (CCR) of 5.77 ppm/K, whereas the contemporaneous temperature response (CTR) is statistically indistinguishable from zero. Accounting for lagged feedbacks, the cumulative temperature response within one millennium of CO2 doubling reaches 11.8 K.
@@ Introduction
The debate over the phase relationship between these variables has been long-standing. Caillon et al. (2003) famously reported that Antarctic temperature leads CO2 by approximately 800 years during Termination III, a finding often used to argue that the initial climate shift is driven by orbital forcing rather than greenhouse gases. However, as Lemoine (2010) pointed out, such simple lead-lag analyses based on peak-to-peak timing fail to disentangle the mutual causal effects from autocorrelation and reverse causality. In a complex system where temperature and carbon are mutually endogenous, a lead in one variable does not preclude the existence of a powerful, simultaneous feedback from the other. Without a framework that accounts for this reciprocity, the magnitude of the positive climate–carbon feedback may be significantly underestimated.
Furthermore, the statistical analysis of paleoclimate records faces fundamental analytical challenges. A primary concern is the high persistence of temperature and CO2 series, which carries the risk of spurious regression (Kaufmann et al., 2006; Cummins et al., 2022). Because these variables exhibit non-stationary, $I(1)$ properties, a cointegration framework is essential to separate true physical bonds from stochastic trends. In this context, Kaufmann and Juselius (2013) provided a seminal benchmark by estimating a Cointegrated Vector Autoregression (CVAR) model; however, as noted by Miller (2019), such models often rely on interpolated data, which can introduce significant statistical artifacts. Furthermore, even when multivariate models are employed to capture slow internal dynamics (Kaufmann and Juselius, 2016), the identification of the contemporaneous interaction matrix ($A_{0}$) has historically relied on pre-assumed physical zero-constraints. Consequently, the objective allocation of causality within a single millennial time step—the very scale where the initial feedback triggers occur—remains an unresolved empirical question.
@@
In this paper, we address these challenges through two methodological advancements. First, we focus our primary structural analysis on a specific 130,000-year interval since the Last Interglacial that is entirely free of observational gaps at a 1,000-year resolution. By utilizing this gap-free sample, we eliminate the need for interpolation, thereby neutralizing the potential biases warned of by Miller (2019). Second, we employ a structural identification strategy that utilizes the strict exogeneity of Milankovitch cycles as an instrumental variable (IV). This allows us to statistically disentangle the contemporaneous temperature response (CTR, $a_{12}$) from the contemporaneous carbon response (CCR, $a_{21}$) without relying on pre-assumed physical constraints on the $A_{0}$ matrix.
@@
Our results reveal a stark asymmetry in the contemporaneous climate–carbon interaction. We find a robust, statistically significant CCR ($\hat{a}_{21}>0$), consistent with the magnitude suggested by last-millennium reconstructions (Frank et al., 2010) and the long-term sensitivity found by Kaufmann and Juselius (2013). In contrast, the CTR ($a_{12}$) is found to be statistically indistinguishable from zero at the millennial scale. This identification implies that while temperature changes trigger an immediate carbon release, the radiative feedback of CO2 onto temperature requires structural lags to manifest. By bridging the gap between time-series econometrics and paleoclimatic physics, we provide a unified explanation for the perceived “synchronicity” in deglacial records as a manifestation of the CCR, while reconciling the “lag” observed in other contexts as the physical adjustment time of the thermal system.
@@
To circumvent this risk while preserving the critical information contained in the levels of the variables, we employ a Vector Error Correction Model (VECM). This framework is appropriate when variables are cointegrated, meaning that while individual series may drift non-stationarily, a stable linear combination exists. In this study, we consider a three-variable system $y_{t}=(T_{t},\ln C_{t},M_{t})^{\prime}$, where $T_{t}$ is Antarctic temperature, $\ln C_{t}$ is the logarithm of CO2 concentration, and $M_{t}$ represents external orbital (Milankovitch) forcing.
We begin with a Structural Vector Autoregressive (SVAR) model that explicitly captures contemporaneous interactions ($A_{0}$):
$y_{t}=A_{0}y_{t}+A_{1}y_{t-1}+\cdots+A_{p}y_{t-p}+u_{t},$ (1)
where $u_{t}\sim\text{i.i.d.}~\mathcal{N}(0,\Sigma_{u})$. The corresponding reduced-form representation is given in Equation (6) in the Appendix. Comparing the two representations, we obtain $A_{i}=(I-A_{0})\Phi_{i}$ for $i=1,\dots,p$, and $u_{t}=(I-A_{0})\epsilon_{t}$.
Moreover, the reduced-form VAR (6) can be reformulated as a VECM:
$\Delta y_{t}=\alpha\beta^{\prime}y_{t-1}+\Gamma_{1}\Delta y_{t-1}+\dots+\Gamma_{p-1}\Delta y_{t-p+1}+\epsilon_{t},$ (2)
where $\beta$ contains the cointegrating vector, $\alpha$ represents the adjustment speeds toward equilibrium, and $\Gamma_{i}$ captures the short-run lagged dynamics. The error term $\epsilon_{t}$ denotes the reduced-form innovations, which are generally correlated element-wise.
@@
To isolate structural shocks ($u_{t}$) from the reduced-form residuals ($\epsilon_{t}$), we utilize the relationship $u_{t}=(I-A_{0})\epsilon_{t}$. This approach follows the framework established by Sims (1980), which provides a basis for analyzing endogeneity and structural innovations within a system of interrelated time series by utilizing the residuals of the estimated model. In our three-variable system, the contemporaneous interaction matrix $A_{0}$ is defined as:
@@
where $s_{ij}$ denotes the sample covariance between $\epsilon_{it}$ and $\epsilon_{jt}$. Equation (3) demonstrates that the allocation of the total contemporaneous correlation between temperature and CO2 into directional causal links is uniquely determined by the sample covariance structure ($s_{ij}$) of the residuals.
@@
The presence of cointegration ensures that temporary structural shocks lead to a permanent shift in the climatic baseline. The long-run impact matrix, defined as $\Xi=\lim_{h\to\infty}\Psi_{h}$ (see Appendix for details), provides a formal bridge between millennial-scale innovations and the long-term equilibrium states.
@@
To further rigorously test the statistical significance of the CTR ($a_{12}$) while accounting for the uncertainty in the CCR ($a_{21}$), we perform a joint estimation using the Generalized Method of Moments (GMM). This approach incorporates both the regression model in Equation (4) and the orthogonality condition derived from the structural shock properties, i.e., the CAC, or Equation (3). The moment conditions for this joint estimation are defined by the structural identification framework established in Section 2.2.
@@
Figure 3: Identification of structural parameters on the Causality Allocation Curve (CAC). The blue line represents the CAC derived from the residual covariance. The red point and dashed ellipse indicate the GMM joint estimate (of Equations (3) and (4)) and the 95% confidence region. The intersection near the $a_{12}=0$ line confirms the unidirectional contemporaneous feedback from temperature to carbon.
@@
Based on the identification results in the previous section, we now evaluate the structural sensitivity of the climate–carbon system. First, to interpret the estimated reduced-form parameters ($\alpha,\beta,\Gamma$) in a structural context, we must consider the transformation $(I-A_{0})$. While the cointegrating vector $\beta$ and the adjustment speeds $\alpha$ remain invariant to this transformation due to the property $\beta^{\prime}\Xi=0$ (as formally derived in Appendix), the short-run lag coefficients $\Gamma$ must be structurally re-identified using the contemporaneous parameters $a_{12}$ and $a_{21}$.
By recalling the structural identity $(I-A_{0})\Delta y_{t}=(I-A_{0})\alpha\beta^{\prime}y_{t-1}+\Gamma^{*}_{1}\Delta y_{t-1}+\dots+u_{t}$, the relationship between the structural and reduced-form coefficients is defined by the identity $\Gamma^{*}_{1}=(I-A_{0})\Gamma_{1}$. The specific element for the $\ln C_{t-1}$ to $T_{t}$ effect can be expanded as follows:
@@
Our structural estimation yields an Antarctic Earth System Sensitivity (ESS) of 13.0 K, which is remarkably consistent with the 11.1 K reported by Kaufmann and Juselius (2013). This alignment is noteworthy given the methodological differences: while Kaufmann and Juselius (2013) utilized interpolated data and physical zero-constraints for identification, our results are derived from a gap-free sample with a data-driven GMM strategy. Specifically, the estimated long-term equilibrium for the 130 ka sample is given by $T-18.748\ln C=\text{const}$. This implies that a doubling of CO2 is associated with a long-term Antarctic increase of $18.748\times\ln(2)\approx 13.0$ K. Applying a polar amplification factor of approximately 0.5, this corresponds to a global ESS of $\sim$6.5 K, aligning with the reconstructed range of $9\pm 4.5$ K by Snyder (2016).
Beyond the long-term equilibrium, the structural VECM reveals the specific causal architecture behind the “slow adjustment” of the climate system. While Kaufmann and Juselius (2016) identified adjustment scales of approximately 19 kyr, our identification of $a_{12}=0$ at the millennial scale explains this delay as a fundamental causal asymmetry. Under our null-CTR identification, the cumulative temperature response within one millennium of a CO2 doubling reaches 11.8 K in Antarctica (or 5.9 K globally).
This response is driven by two mechanisms: the structural lag effect $(\Gamma^{*}_{1})_{12}\times\ln(2)\approx 10.5$ K and the error correction process $\alpha\times(-13.0)\approx 1.3$ K. This result implies that while temperature triggers an immediate carbon release (CCR), the reciprocal radiative impact of CO2 onto temperature is not instantaneous but manifests through structural lags.
Finally, we compare these paleo-derived responses with modern climate observations. A 1,000-year response of 5.9 K global warming may appear high compared to the 1.2 K increase observed since the industrial revolution. However, this discrepancy is consistent with the non-linear “front-loading” of climate response: the immediate warming (Transient Climate Response) is followed by centuries of “committed” warming from slow internal feedbacks. The gap between the modern 1.2 K and our 1,000-year projection quantifies the massive “thermal debt” inherent in the Earth system’s long-term dynamics.
Our GMM estimate $\hat{a}_{21}=0.0206$ corresponds to a carbon cycle sensitivity of $\sim$5.77 ppm/K, which is highly consistent with the $7.7$ ppm/K reported by Frank et al. (2010). The stability of this CCR across different frameworks reinforces the warning by Kaufmann and Juselius (2016) that univariate analyses are prone to bias. By explicitly modeling the joint endogeneity, we ensure that our sensitivity parameters reflect the integrated dynamics of the climate–carbon system rather than isolated correlations.
@@
Figure 4 illustrates the response to the CO2 doubling shock. Consistent with our null-CTR identification, temperature shows no immediate response within the same 1,000-year step. Instead, the warming is triggered by structural lag effects, gradually converging toward the long-term ESS of 13.0 K in Antarctica. This dynamic visualization aligns with Rohling et al. (2012), who emphasize that climate sensitivity is a function of the timescale, with the full magnitude of slow feedbacks becoming apparent only as the system approaches a new steady state.
@@
Appendix: The VECM Representation and Impulse Responses
@@
Assuming a constant shock has been applied from the past to the present ($\epsilon_{t-h}=\epsilon$), Equation (7) yields $y_{t}=(\sum_{h=0}^{\infty}\Psi_{h})\epsilon$. On the other hand, assuming a long-run equilibrium state ($y_{t}=y_{t-1}=\dots=y$) and substituting this into Equation (6), we obtain $\left(I-\sum_{i=1}^{p}\Phi_{i}\right)y=Gy=\epsilon$. By comparing the two, the following relationship is derived:
@@
where $C_{j}=\Psi_{j}-\Psi_{j-1}$ represents the impact of the shock $\epsilon_{t-j}$ on the current “change” $\Delta y_{t}$. Given that the level $y_{t}$ is the cumulative sum of past differences $y_{t}=y_{0}+\sum_{j=1}^{t}\Delta y_{j}$, the impulse response for the level, $\Psi_{h}=C_{h}+\Psi_{h-1}$, can be rewritten as:
@@
$\Psi_{h}=\sum_{j=0}^{h}C_{j}=\sum_{j=0}^{\infty}C_{j}-\sum_{j=h+1}^{\infty}C_{j}=\Xi+\Upsilon_{h}$
@@
Note that $\Phi_{1}=I+\Pi+\Gamma_{1}$, $\Phi_{i}=\Gamma_{i}-\Gamma_{i-1}$, and $\Phi_{p}=-\Gamma_{p-1}$. Given the non-stationarity of the level series $y_{t-1}$, $\Pi$ cannot be of full-rank. The rank-dificient case $\Pi=\alpha\beta^{\prime}$ represents the long-run equilibrium (cointegrating) relationships. Engle and Granger (1987) demonstrated that for an $I(1)$ series $y_{t}$, a linear combination $\beta^{\prime}y_{t}$ exists that is stationary, since $\Delta y_{t}$ in Equation (11) must all be $I(0)$.
@@
This property carries a profound physical meaning: while an exogenous shock (such as a sudden change in orbital forcing or a carbon cycle innovation) can permanently shift the absolute levels of temperature and CO2, it is fundamentally incapable of breaking the equilibrium “bond” between them. The variables move to a new climatic baseline, but the underlying physical relationship, defined by $\beta$, remains intact.